OpenStudy (anonymous):

What is the vertex of the parabola given by the equation y = 9(x + 4)^2 – 3?

OpenStudy (anonymous):

the base equation is y=a(x-h)^2 + k vertex of parabola is as (h,k) what is your vertex of parabola? (:

OpenStudy (anonymous):

would it be (-4, -3)? since the formula has a negative, you have to negate it right?

OpenStudy (whpalmer4):

If you have the parabola in form \[y = ax^2+bx+c\] which in this case would be \[y = 9x^2 +72x + 141\]the x-value of the vertex is given by \[x = -\frac{b}{2a} = -\frac{72}{2*9} = \] The y-value you find by substituting that value of \(x\) into the equation.

OpenStudy (whpalmer4):

yes, that is correct.

OpenStudy (anonymous):

ah the negative. there's a foolproof way of finding out: 0=a(x-h)^2 make x your guessed answer, which is -4: 2((-4)+4)^2=? if you get zero, then its your answer. if you don't just try the opposite ^^

OpenStudy (whpalmer4):

Here's a possibly slightly easier to understand way to find out: plug the point into the equation. Does it work? The vertex, after all, is a point on the parabola. Let's say we guessed wrong, and said (4,-3): \[y = 9(x+4)^2-3\]\[-3=9(4+4)^2-3\]\[-3=9(64)-3\]Uh, that's clearly not the right one. \[-3 = 9(-4+4)^2-3\]\[-3=9(0)-3\]Now we're talking...

OpenStudy (anonymous):

((-4)+4)^2=0 works too actually you don't even have to add the a or 9 :D

OpenStudy (anonymous):

Auxuris, thank you for giving me formulas to find the vertex with. Whpalmer4, thank you for explaining an easy way to see if the point fits. Thank you both so much! This helped me alot! I really wish I could give both of you a medal :(